OpenStudy (anonymous):

there are eight job applicants sitting in a waiting room - four women and four men. If two of the applicants are selected at random, what is the probability that both will be men?

ganeshie8 (ganeshie8):

how many total ways u can choose "two" applicants from "eight" applicants ?

OpenStudy (anonymous):

28??

OpenStudy (anonymous):

16?

ganeshie8 (ganeshie8):

u can choose "two" applicants from "eight" applicants in \(\large ^8C_2 = 28\) ways

OpenStudy (anonymous):

how can we calculate without using calculator??

ganeshie8 (ganeshie8):

and u can choose "two" men from a group of "four" men in \(\large ^4C_2 = 6\) ways

OpenStudy (anonymous):

but i'm not allowed to use calculator

ganeshie8 (ganeshie8):

so the probability for choosing both men = \(\large \frac{6}{28} = \frac{3}{14}\)

ganeshie8 (ganeshie8):

you dont need calculator

ganeshie8 (ganeshie8):

use below formula :- \(\large ^nC_r = \frac{n!}{r!(n-r)!}\)

OpenStudy (anonymous):

oh okayy ... i know this formula.. thanks alot!!

ganeshie8 (ganeshie8):

\(\large ^8C_2 = \frac{8!}{2!(8-2)!} =\frac{8!}{2!6!} = \frac{8\times 7\times 6!}{2!6!} = \frac{8\times 7}{2} = 28\)

ganeshie8 (ganeshie8):

np... u wlc :)

OpenStudy (anonymous):

i have few more probabilities question... will ya help??

ganeshie8 (ganeshie8):

sure il try :)